# STBA: Towards Evaluating the Robustness of DNNs for Query-Limited Black-box Scenario

### v1

@@ Abstract
Figure 1: The clean example and their corresponding adversarial example generated by baselines and the proposed STBA. The first one is the clean images, and the flowings are the adversarial example generated by Square Attack [1], RS_Attack [2], ASH [3] and NPAttack [4], respectively.
@@ I Introduction
Deep neural networks (DNNs) are susceptible to adversarial examples [5, 6], which are crafted by subtly perturbing a clean input [7, 8, 9], especially for computer vision (CV) tasks, such as image classification, face recognition [10], and object detection. Except for the threat brought by such attacks, adversarial examples are the most effective way to evaluate existing models’ vulnerability and can further impel the defender to improve the robustness of existing models and guide the design of new safety models. The critical point in carrying out adversarial attacks on CV models is how to generate adversarial examples with high imperceptibility and attack success rate. Various methods have been proposed to build adversarial examples; among them, the black-box attacks are more practical because they do not need to access the target model and only require the prediction results [11, 12, 13, 14], which makes them are more suitable to destroy the DNN models deployed in the physical world.
Although most existing black-box attacks can obtain a high success rate by adding noise to the original image, they are not ideal in terms of naturalness since the added perturbations are not harmonious with the original clean image [15]. Meanwhile, most existing black-box attacks are query and optimization-based, which regard the process of calculating adversarial examples (or adversarial noise) as an optimization problem. They iteratively optimize the intermediate adversarial example with the query results until it can attack the target model successfully [16]. Those methods require a vast number of queries to ensure that most crafted adversarial examples for a dataset can be used to attack victim models successfully, which results in a heavy computation and time burden; worse, a mass of queries can be easily detected by defenders. Finally, most deployed DNN models are enhanced by various defense mechanisms; among them, adversarial training, which finetunes the DNN model on the pre-collected adversarial examples, is verified as the most effective way to improve the model’s robustness against unknown attacks. Since such adversarial examples are often locally generated with noise-adding attacks, existing black-box attacks are more challenging to successfully attack these robust models.
@@
To better evaluate the vulnerability and robustness of existing DNN models, including the robust ones, and improve the concealment of adversarial examples and the query efficiency of black-box attacks under the query-limited scenario, we formulate the issue of synthesizing adversarial examples beyond additive perturbations and propose a novel non-addition black-box attack method called STBA. More specifically, STBA uses spatial transformation techniques to generate adversarial examples with better concealment rather than directly adding well-designed noise to the benign image. The spatial transform technique can generate an eligible adversarial example by leveraging a well-optimized smooth flow field matrix $\bm{f}$ to move each pixel to a new location. Besides, to further enhance the imperceptibility of the generated adversarial noise, we first split the high- and low-frequency parts from clean images and then only apply the will-estimated flow field $\bm{f}$ to the high-frequency part to formulate the adversarial high-frequency counterpart, which will be combined to the original low-frequency part to synthesis the final adversarial examples. Furthermore, to address the query challenge of optimization-based black attacks, STBA leverages the scarce information, query accounts, attack success rate, and the risk of exposure in the query-limited situation, employed with gradient estimation to optimize such a flow field, which can effectively find the global optima with higher probability but require less information from the target model and straightforward the generation process. Therefore, the proposed STBA can achieve a higher attack success rate with a harsher query budget.
We draw the adversarial examples of different methods in Fig. 1, where the victim model is ResNet-152 [23] and the image is from the ImageNet [24] dataset. The visual results show that our proposed method alters the clean image slightly, thus ensuring the invisibility of the adversarial perturbations. Furthermore, we conducted extensive experiments on four different CV benchmark datasets; the statistical results indicate that the STBA can obtain the highest attack success rate with the max query budget limited as $Q_{max}=1000$ or $Q_{max}=10000$. Specifically, STBA can achieve a nearly 100% attack success rate on the normally trained models when $Q_{max}=1000$ and on the robust trained models when $Q_{max}=10000$ in most cases, while the baselines are lower than 90% and 70% in the most cases, respectively. Furthermore, we found that STBA can achieve the highest attack success rate with fewer average queries compared to the existing query-limited black-box methods on the target victim models while outperforming them concerning attack success rate. The main contributions could be summarized as follows:
@@
  • For further enhancing the imperceptibility of adversarial examples, we integrate spatial transform and Gaussian blur techniques into our attack framework. Specifically, we first decomposed the image into high-frequency and low-frequency components and then only applied the calculated spatial transform to the high-frequency part to improve the generated adversarial images’ visibility and image quality.
@@ II-A Spatial transform-based attacks
The concept of spatial transformation is first mentioned in [25], which indicates that conventional neural networks are not robust to rotation, translation, and dilation. Next, [26] utilized the spatial transform techniques and proposed the stAdv method to craft adversarial examples with a high fooling rate and perceptually realistic beyond noise-adding way. They change each pixel position in the clean image by applying a well-designed flow field matrix $\bm{f}$ to it clean image to formulate the adversarial counterpart. Later, [27] proposed a new method to produce the universal adversarial examples by combining the spatial transform and pixel-level distortion, and it successfully increased the attack success rate against universal perturbation to more than 90%. In the literature [28], the authors applied spatial transform to the YUV space to generate adversarial examples with higher superiority in image quality. Also, Liu et al. [29] utilized spatial transformations by applying the flow field to the latent representations from the normalized flow model [30] to guarantee that generated adversarial examples are more perceptually similar to the original image.
@@ II-B Query-Efficient black-box attack
Although most optimized-based black-box attacks have achieved a high attack success rate, even reaching 100% in some cases, the tremendous need for queries to optimize adversarial examples is still a big challenge. Empirically, the average number of queries required to obtain a suitable adversarial example is from thousands to hundreds of thousands [22, 31, 32], which is obviously not feasible in the physical world and can be detected by the target DNN models easily. Therefore, recently, researchers have proposed various methods to generate adversarial examples under query-limited queries or in a query-efficient way [33, 34].
@@
The first work to limit the number of queries in black-box attacks is the Bayesian Optimization-Based Black-box Attack [35]. It uses the Bayesian optimization strategy to reduce the average number of queries to 1/10 of the random perturbation strategies. Then, Ilyas et al. proposed NES [31] to improve the query efficiency while maintaining a higher success rate in query and information-limited settings. AdvFlow [19] improves the attack performance with a limited query budget by disturbing the image’s latent representation obtained from pre-trained Normalize Flow. Square Attack [1] uses square-like adversarial perturbation to extend attacks with random search. Accelerated Sign Hunter (ASH) [3] crafts the adversarial example by applying a Branch Prune Strategy that infers the unknown sign bits to avoid unnecessary queries and further improve the query efficiency within a limited query budget.
@@
Besides, some researchers have proposed combining the transferability of adversarial attacks to improve query efficiency, which is dubbed query- and transfer-based attacks. F. Suya et al. [36] combined transfer-based attack [37] and optimize-based black-box and proposed Hybrid Batch Attacks (HBA). The HBA can use the information of the existing local models to generate candidate adversarial seeds and use these seeds to query the target model and optimize the generated images iteratively. Its seed-first strategy saves about 70% query times for an attack. In [38], the authors propose a query-efficient boundary-based attack (QEBA) that only relies on the predicted hard labels and presents an optimal analysis of gradient estimation based on dimensional reduction and gets good results. In [39], Dong et al. proposed two prior-guided gradient-free stochastic algorithms (PRGF) based on biased sampling and gradient averaging, respectively, which require fewer queries while guaranteeing a higher attack success rate. CG-Attack [40] combines a well-trained c-glow model and CAM-ES to extend attacks. MCG-Attack [41] improves the query efficiency by fully utilizing both example-level and model-level transferability.
@@
Figure 2: Overview of Spatial Black-box Transform Attack. The $\bm{x}$ and $\bm{x}^{adv}$ are the benign image and the corresponding adversarial counterpart, respectively. The ${\bm{x}_{high}}$ and $\bm{x}_{low}$ represent the high-frequency part and the low-frequency part of the benign image. The $\otimes$ represents the spatial transformation operation, and $\oplus$ means element addition. The $\bm{f}_{i},i\in\{1,2,...,n\}$ are the mini-batch sampled candidate flow field sampled from an isometric normal distribution $\mathcal{N}$, and the $\bm{f}^{adv}$ is the final optimized flow field that applies to the original clean image to formulate the final adversarial image $\bm{x}^{adv}$.
@@
However, these transfer-based attacks heavily rely on the similarity between the substitute model and the oracle model. In this paper, we only focus on score-based attacks, which are independent of local surrogate models. Therefore, for adversarial attacks, especially for the Query-Efficient black-box attacks, pose the request for a method that is efficient and effective in building adversarial examples with high imperceptibility and quality in performing attacks on unknown DNN models within limited queries. On the other hand, with the development of defense mechanisms, higher requirements are placed on the invisibility of adversarial examples. To achieve these goals, we learn from the previous studies that adversarial examples can be gained in a beyond noise-adding way. Hence, we are well motivated to develop a novel method to disturb the original image by applying a well-calculated flow field $\bm{f}$ to clean images to generate adversarial ones, which makes it possible to build qualified adversarial examples with high invisibility in dozens or hundreds of queries.
@@ III-A Problem Definition
Different from this, we introduce the spatial transform techniques to high-frequency from the clean image to build the adversarial example $\bm{x}^{adv}$. As illustrated in Fig. 2, the proposed Spatial Transform Black-box Attack framework can be divided into flowing stages: The first stage is decomposing the different frequency information, i.e., the high-frequency part ${\bm{x}_{high}}$ and the low-frequency part $x_{low}$, from the original clean image $\bm{x}$ with the help of Gaussian Blur. Second, we random sampling a mini-batch candidate flow field $\bm{f}_{i},i\in\{1,2,3...,n\}$ from an isometric normal distribution $\mathcal{N}$, which will be applied to the high-frequency part ${\bm{x}_{high}}$, and then will be combined with the low-frequency part $\bm{x}_{low}$ to build the candidate adversarial example used to query the black-box victim model to estimate the gradient in an iterative manner. Once the gradient is estimated, it can used to guide the optimization process to update the flow field $\bm{f}^{adv}$. Finally, we can obtain the adversarial example $\bm{x}^{adv}$ by applying the calculated flow field $\bm{f}^{adv}$ to the high-frequency part ${\bm{x}_{high}}$.
@@ III-B Adversarial Example Generation
We utilize the spatial transform method to build adversarial examples in a non-noise additional way. The spatial transform techniques using a flow field matrix $\bm{f}\in[0,1]^{2*h*w}$ to transform the original image $\bm{x}$ to $\bm{x}_{st}$ [26], where $h$ and $w$ are the height and width of the image. In particular, for our proposed STBA, we only need to apply such transform operation to the high-frequency part, i.e., ${\bm{x}_{high}}$. Specifically, assume the input is ${\bm{x}_{high}}$ and its transformed counterpart $\bm{x}_{high}^{st}$, for the $i$-th pixel in $\bm{x}_{high}^{st}$ at the pixel location $(u_{st}^{i},v_{st}^{i})$, we need to calculate the flow field $\bm{f}_{i}=(\Delta u^{i},\Delta v^{i})$. So, the $i$-th pixel $\bm{x}_{high}^{i}$’s location in the transformed image can be indicated as:
@@
${\bm{x}_{high}^{st}}^{i}=\sum_{q\in\bm{N}(u^{i},v^{i})}\bm{x}^{q}(1-|u^{i}-u^{q}|)(1-|v^{i}-v^{q}|),$ (4)
where $\bm{N}(u^{i},v^{i})$ is the neighborhood, that is, the four positions (top-left, top-right, bottom-left, bottom-right) tightly surrounding the target pixel $(u^{i},v^{i})$. For each pixel ${\bm{x}_{high}^{st}}^{i}$ in the adversarial high-frequency part $\bm{x}_{high}$, we can calculate it with with Eq. 4 to formulate the $\bm{x}_{high}^{st}$, i.e., the final adversarial high-frequency part $\bm{x}_{high}^{adv}$ for adversarial attack. Once the $\bm{f}$ has been computed, we can obtain the $\bm{x}^{adv}$ by applying such flow field $\bm{f}$, which is given by:
$\bm{x}_{high}^{adv}=\sum_{q\in\bm{N}(u^{i},v^{i})}\bm{x}_{high}^{q}(1-|u^{i}-u^{q}|)(1-|v^{i}-v^{q}|)),$
@@ III-C Objective Functions
For adversarial attacks, the goal is making $\bm{\mathcal{C}}(\bm{x}^{adv})\neq y$. We give the objective function as:
@@
where $\lambda$ is a hyper-parameter to balance each loss item, in this paper, we set it as $\lambda=5$ (The reason why we set $\lambda=5$, please refer to Sec. IV-F ).
Algorithm 1 Spatial Transform for Query-Efficient Attack
Parameter: The flow field $\bm{f}=[0,1]^{2*h*w}$.
Output: The adversarial example $\bm{x}^{adv}$ used for attack.
1:  Initialize the flow budget $\xi$ with $\xi=10^{-1}$.
2:  Initialize $\bm{\mu}$ randomly.
3:  Initialize the parameters of the flow $\bm{f_{0}}$ with zeros.
4:  Decompose $\bm{x}$ to $\bm{x}_{high}$ and $\bm{x}_{low}$
5:  while $q<=\bm{Q}_{max}/n_{sample}$ do
7:     Set $\bm{f}_{k}=\bm{\delta_{f_{k}}}$ for all $k=1,...,n_{sample}$.
8:     Compute ${\bm{x}_{high}^{adv}}^{k}$ by Eq. 14 for all $k=1,...,n_{sample}$.
[…]
11:     Compute $\nabla_{\bm{\mu}}\bm{J}(\bm{\mu},\sigma)=\frac{1}{n_{sample}}{\textstyle\sum_{k=1}^{n_{sample}}}\hat{\mathcal{L}}\cdot\bm{\epsilon}_{k}$
12:     update $\bm{\mu}\longleftarrow\bm{\mu}-lr\nabla_{\bm{\mu}}\bm{J}(\bm{\mu},\sigma)$
13:     if $q\ \%\ adjust_{num}=0$ then
14:        Update $\xi_{t}$ by Eq.8.
15:     end if
17:     if $\bm{x}^{adv}_{t}$ attack $\bm{\mathcal{C}}$ successfully then
18:        break.
19:     end if
20:  end while
21:  return $\bm{x}^{adv}_{t}$

### v2

@@ Abstract
Figure 1: The clean example and their corresponding adversarial example generated by baselines and the proposed STBA. The first one is the clean images, and the flowings are the adversarial example generated by Square Attack [1], AdvFlow [2], RS_Attack [3], ASH [4], NPAttack [5], and DifAttack [6], respectively.
@@ I Introduction
Deep neural networks (DNNs) are susceptible to adversarial examples [7, 8, 9], which are crafted by subtly perturbing a clean input [10, 11, 12, 13], especially for computer vision (CV) [14] and Natural Language Processing (NLP) tasks, such as image classification, face recognition [15], and language translation [16]. Except for the threat brought by such attacks, adversarial examples are the most effective way to evaluate existing models’ vulnerability and can further impel the defender to improve the robustness of existing models and guide the design of new safety models. The critical point in carrying out adversarial attacks on CV models is how to generate adversarial examples with high imperceptibility and attack success rate. Various methods have been proposed to build adversarial examples; among them, the black-box attacks are more practical because they do not need to access the target model and only require the prediction results [17, 18, 19, 20], which makes them are more suitable to destroy the DNN models deployed in the physical world.
Although most existing black-box attacks can obtain a high success rate by adding noise to the original image, they are not ideal in terms of naturalness since the added perturbations are not harmonious with the original clean image [21]. Meanwhile, most existing black-box attacks are query and optimization-based, which regard the process of calculating adversarial examples (or adversarial noise) as an optimization problem. They iteratively optimize the intermediate adversarial example with the query results until it can attack the target model successfully [22]. Those methods require a vast number of queries to ensure that most crafted adversarial examples for a dataset can be used to attack victim models successfully, which results in a heavy computation and time burden; worse, a mass of queries can be easily detected by defenders. Finally, most deployed DNN models are enhanced by various defense mechanisms; among them, adversarial training, which finetunes the DNN model on the pre-collected adversarial examples, is verified as the most effective way to improve the model’s robustness against unknown attacks. Since such adversarial examples are often locally generated with noise-adding attacks, existing black-box attacks are more challenging to attack these robust models successfully.
@@
To better evaluate the vulnerability and robustness of existing DNN models, including the robust ones, and improve the concealment of adversarial examples and the query efficiency of black-box attacks under the query-limited scenario, we formulate the issue of synthesizing adversarial examples beyond additive perturbations and propose a novel non-addition black-box attack method called STBA. More specifically, STBA uses spatial transformation techniques to generate adversarial examples with better concealment rather than directly adding well-designed noise to the benign image. The spatial transform technique can generate an eligible adversarial example by leveraging a well-optimized smooth flow field matrix $\bm{f}$ to move each pixel to a new location. Besides, to further enhance the imperceptibility of the generated adversarial noise, we first split the high- and low-frequency parts from clean images and then only apply the will-estimated flow field $\bm{f}$ to the high-frequency part to formulate the adversarial high-frequency counterpart, which will be combined to the original low-frequency part to synthesis the final adversarial examples. Furthermore, to address the query challenge of optimization-based black-box attacks, STBA leverages the scarce information, query accounts, attack success rate, and the risk of exposure in the query-limited situation, employed with gradient estimation to optimize such a flow field, which can effectively find the global optima with higher probability but require less information from the target model and straightforward the generation process. Therefore, the proposed STBA can achieve a higher attack success rate with a harsher query budget.
We draw the adversarial examples of different methods in Fig. 1, where the victim model is ResNet-152 [28] and the image is from the ImageNet [29] dataset. The visual results show that our proposed method alters the clean image slightly, thus ensuring the invisibility of the adversarial perturbations. Furthermore, we conducted extensive experiments on four different CV benchmark datasets, the statistical results indicate that the STBA can obtain the highest attack success rate with the max query budget limited as $Q_{max}=1000$ or $Q_{max}=10000$. Specifically, STBA can achieve a nearly 100% attack success rate on the normally trained models when $Q_{max}=1000$ and on the robust trained models when $Q_{max}=10000$ in most cases, while the baselines are lower than 90% and 70% in the most cases, respectively. Furthermore, we found that STBA can achieve the highest attack success rate with fewer average queries compared to the existing query-limited black-box methods on the target victim models while outperforming them concerning attack success rate. The main contributions could be summarized as follows:
@@
  • To further enhance the imperceptibility of adversarial examples, we integrate Gaussian blur techniques and dynamical strategy into our attack framework. Specifically, we first decomposed the image into high-frequency and low-frequency components. We then only applied the dynamically constrained flow field $\bm{f}$ to the high-frequency part to improve the generated adversarial images’ visibility and image quality.
@@ II-A Spatial transform-based attacks
The concept of spatial transformation is first mentioned in [30], which indicates that conventional neural networks are not robust to rotation, translation, and dilation. Next, [31] utilized the spatial transform techniques and proposed the stAdv method to craft adversarial examples with a high fooling rate and perceptually realistic beyond noise-adding way. They change each pixel position in the clean image by applying a well-designed flow field matrix $\bm{f}$ to the clean image to formulate the adversarial counterpart. Later, [32] proposed a new method to produce the universal adversarial examples by combining the spatial transform and pixel-level distortion, and it successfully increased the attack success rate against universal perturbation to more than 90%. In the literature [33], the authors applied spatial transform to the YUV space to generate adversarial examples with higher superiority in image quality. Also, Liu et al. [34] utilized spatial transformations by applying the flow field to the latent representations from the normalized flow model [35] to guarantee that generated adversarial examples are more perceptually similar to the original image.
@@ II-B Query-Efficient black-box attack
Although most optimized-based black-box attacks have achieved a high attack success rate, even reaching 100% in some cases, the tremendous need for queries to optimize adversarial examples is still a big challenge. Empirically, the average number of queries required to obtain a suitable adversarial example is from thousands to hundreds of thousands [27, 36], which is obviously not feasible in the physical world and can be detected by the target DNN models easily. Therefore, recently, researchers have proposed various methods to generate adversarial examples under query-limited queries or in a query-efficient way [37, 38].
@@
The first work to limit the number of queries in black-box attacks is the Bayesian Optimization-Based Black-box Attack [39]. It uses the Bayesian optimization strategy to reduce the average number of queries to 1/10 of the random perturbation strategies. Then, Ilyas et al. proposed NES [36] to improve the query efficiency while maintaining a higher success rate in query and information-limited settings. In [40], the authors propose a query-efficient boundary-based attack (QEBA) that only relies on the predicted hard labels and presents an optimal analysis of gradient estimation based on dimensional reduction and gets good results. AdvFlow [2] improves the attack performance with a limited query budget by disturbing the image’s latent representation obtained from pre-trained Normalize Flow. Square Attack [1] uses the random searched square-like adversarial perturbation to extend attacks. Accelerated Sign Hunter (ASH) [4] crafts the adversarial example by applying a Branch Prune Strategy that infers the unknown sign bits to avoid unnecessary queries and further improve the query efficiency within a limited query budget.
@@
Besides, some researchers have proposed combining the transferability of adversarial attacks to improve query efficiency, which is dubbed query- and transfer-based attacks. F. Suya et al. [41] combined transfer-based attack [42] and optimize-based black-box and proposed Hybrid Batch Attacks (HBA). The HBA can use the information of the existing local models to generate candidate adversarial seeds and use these seeds to query the target model and optimize the generated images iteratively. Its seed-first strategy saves about 70% query times for an attack. MCG-Attack [43] improves the query efficiency by fully utilizing both example-level and model-level transferability. DifAttack [6] disentangles image latent into adversarial and visual features, trains an autoencoder with clean images and their adversarial examples of surrogate model(s), and iteratively optimizes adversarial features based on victim model feedback to create successful AEs while maintaining visual consistency.
@@
Figure 2: Overview of Spatial Black-box Transform Attack (STBA). The Black-box Model is the victim model. The $\bm{x}$ and $\bm{x}^{adv}$ are the benign image and the corresponding adversarial counterpart, respectively. The ${\bm{x}_{high}}$ and $\bm{x}_{low}$ represent the high-frequency part and the low-frequency part of the benign image. The $\otimes$ represents the spatial transformation operation, and $\oplus$ means element addition. The $\bm{f}_{i},i\in\{1,2,...,n\}$ are the mini-batch candidate flow field sampled from an isometric normal distribution $\mathcal{N}$, and the $\bm{f}^{adv}$ is the final optimized flow field that applies to the original clean image to formulate the final adversarial image $\bm{x}^{adv}$.
@@
However, these transfer-based attacks heavily rely on the similarity between the substitute model and the oracle model. In this paper, we only focus on score-based attacks, which are independent of local surrogate models. Therefore, adversarial attacks, especially for the Query-Efficient black-box attacks, pose the request for a method that is efficient and effective in building adversarial examples with high imperceptibility and quality in performing attacks on unknown DNN models within limited queries. On the other hand, with the development of defense mechanisms, higher requirements are placed on the invisibility of adversarial examples. To achieve these goals, we learn from the previous studies that adversarial examples can be gained in a beyond-noise-adding way. Hence, we are well motivated to develop a novel method to disturb the original image by applying a well-calculated flow field $\bm{f}$ to clean images to generate adversarial ones, which makes it possible to build qualified adversarial examples with high invisibility in dozens or hundreds of queries.
@@ III-A Problem Definition
Different from this, we introduce the spatial transform techniques to the high-frequency part of the clean image to build the adversarial example $\bm{x}^{adv}$. As illustrated in Fig. 2, the proposed Spatial Transform Black-box Attack framework can be divided into flowing stages: The first stage is decomposing the different frequency information, i.e., the high-frequency part ${\bm{x}_{high}}$ and the low-frequency part $x_{low}$, from the original clean image $\bm{x}$ with the help of Gaussian Blur (we set the kernel size as 3x3). Second, we random sampling a mini-batch candidate flow fields $\bm{f}_{i},i\in\{1,2,3...,n\}$ from an isometric normal distribution $\mathcal{N}$, which will be applied to the high-frequency part ${\bm{x}_{high}}$, and then combined with the low-frequency part $\bm{x}_{low}$ to build the candidate adversarial examples used to query the black-box victim model to estimate the gradient in an iterative manner. Once the gradient is estimated, it can used to guide the optimization process to update the flow field $\bm{f}^{adv}$. Finally, we can obtain the adversarial example $\bm{x}^{adv}$ by applying the calculated flow field $\bm{f}^{adv}$ to the high-frequency part ${\bm{x}_{high}}$.
@@ III-B Adversarial Example Generation
We utilize the spatial transform method to build adversarial examples in a non-noise additional way. The spatial transform techniques using a flow field matrix $\bm{f}\in[0,1]^{2*h*w}$ to transform the original image $\bm{x}$ to $\bm{x}_{st}$ [31], where $h$ and $w$ are the height and width of the image. In particular, for our proposed STBA, we only need to apply such transform operation to the high-frequency part, i.e., ${\bm{x}_{high}}$. Specifically, assume the input is ${\bm{x}_{high}}$ and its transformed counterpart $\bm{x}_{high}^{st}$, for the $i$-th pixel in $\bm{x}_{high}^{st}$ at the pixel location $(u_{st}^{i},v_{st}^{i})$, we need to calculate each element of the flow field as $\bm{f}_{i}=(\Delta u^{i},\Delta v^{i})$. So, the $i$-th pixel $\bm{x}_{high}^{i}$’s location in the transformed image can be indicated as:
@@
${\bm{x}_{high}^{st}}^{i}=\sum_{q\in\bm{N}(u^{i},v^{i})}\bm{x}_{high}^{q}(1-|u^{i}-u^{q}|)(1-|v^{i}-v^{q}|),$ (4)
where $\bm{N}(u^{i},v^{i})$ is the neighborhood, that is, the four positions (top-left, top-right, bottom-left, bottom-right) tightly surrounding the target pixel $(u^{i},v^{i})$. For each pixel ${\bm{x}_{high}^{st}}^{i}$ in the adversarial high-frequency part $\bm{x}_{high}$, we can calculate it with with Eq. 4 to formulate the $\bm{x}_{high}^{st}$, i.e., the final adversarial high-frequency part $\bm{x}_{high}^{adv}$. Once the $\bm{f}$ for formulating $\bm{x}_{high}^{adv}$ has been computed, we can obtain the final $\bm{x}^{adv}$ by adding such adversarial high-frequency part $\bm{x}_{high}^{adv}$ with low-frequency part $\bm{x}_{low}$ of the clean image, which is given by:
@@ III-C Objective Functions
For adversarial attacks, the goal is making $\bm{\mathcal{C}}(\bm{x}^{adv})\neq y$. We give the objective function as follows:
@@
where $\lambda$ is a hyper-parameter to balance each loss item, in this paper, we set it as $\lambda=5$ (For the reason why we set $\lambda=5$, please refer to Sec. IV-F ).